a ball is dropped on to the floor from a height of 10 m. it rebounds to aheight of 2.5m. if thw ball is in contact with the floor for 0.01seconds. find averge accelaration

Grade:12

2 Answers

Nikhil
10 years ago
Swetha
10 years ago
consider upwards as positive and down wards as negative. it falls from a height of 10 m. so on reaching the ground it has a velocity of squareroot(2*g*10) = 10*1.414 = 14.14m/s in the downward direction. (taking g=10m/s) as it rebounds to a height of 2.5m its velocity soon after leaving the ground is squareroot(2*g*2.5) = 5*1.414 = 7.07 m/s in the upward direction both the above values can be found from the equation v^2 - u^2 = 2*a*s we have average accleration = change in velocity / time taken = 7.07 - (-14.14) / 0.01 = 2121 m/s^2 that is 2121 m/s^2 in the upward direction.